Suppose u, v and w are vectors in 3-space,where u =(u1u2, u3), v =(v1, v2, v3) and w= (w1, w2, w3).

Express ( u * v) * w as a determinant.
1
Expert's answer
2020-04-03T16:58:28-0400

"u=(u_1,u_2,u_3), v=(v_1,v_2,v_3), w=(w_1,w_2,w_3)"

"u\times v=\begin{vmatrix}
 \vec{i} &\vec{j}&\vec{k} \\
 u_1 &u_2&u_3\\
v_1&v_2&v_3
\end{vmatrix}=\\
=\vec{i}\cdot(u_2v_3-u_3v_2)-\vec{j}\cdot(u_1v_3-u_3v_1)+\\
+\vec{k}\cdot(u_1v_2-u_2v_1)=\\
=(u_2v_3-u_3v_2,-u_1v_3+u_3v_1,u_1v_2-u_2v_1)\\
(u\times v)\times w=\\
=\begin{vmatrix}
 \vec{i} &\vec{j}&\vec{k} \\
 u_2v_3-u_3v_2&-u_1v_3+u_3v_1&u_1v_2-u_2v_1\\
w_1&w_2&w_3
\end{vmatrix}"